Find differential equations dy/dx, if  y = sec-1 (1+x2​/2x)

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
it is not called find differential equation.
It is called as find derivatives.
 
Now secy = 1 + x^2 /2x
 
sec y = 1/2x + x/2
 
sec y tan y dy/dx = ln(x)/2 + ½
 
Hence dy/dx = (1 + ln x) /(2 ( sec y tan y))